OpenStudy (anonymous):

Which of the following equations best represents the graph of f(n)? f(y) = n f(x) = n x = f(n) y = f(n)

jimthompson5910 (jim_thompson5910):

y and f(x) are usually interchangeable ie y = f(x) If we replace x with n, we get y = f(n) where n is the input now instead of x.

OpenStudy (anonymous):

Thanks, can you help me with one more?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

The gralow plots a function f(x).If x represent time, the average rate of change of the function f(x) in the first three seconds is ___.

jimthompson5910 (jim_thompson5910):

first you need f(0) and f(3)

jimthompson5910 (jim_thompson5910):

f(0) = ? f(3) = ?

OpenStudy (anonymous):

100 and 200? Because the y intercept is 100 at 0 seconds and 200 at 3 seconds

jimthompson5910 (jim_thompson5910):

See attached

jimthompson5910 (jim_thompson5910):

Notice I added in those blue lines

OpenStudy (anonymous):

Oops. Accident 100 and 220?

jimthompson5910 (jim_thompson5910):

Subtract the y values: 220-100 = 120 Subtract the x values: 3-0 = 3 Then divide the differences: ydiff/xdiff = 120/3 = 40 So the slope of this line is 40

jimthompson5910 (jim_thompson5910):

The slope is the same as the average rate of change (since we're talking about straight lines)